a sodium hydroxide solution contains 40% by weight of NaOH has specific gravity= 1.5. wat volume of this soln. is required to get 500 ml of .45M SOLUTION

Grade:

1 Answers

vikas askiitian expert
13 years ago

40% by weight means , 40gm NaOH in 100 gm solution..

dincity of solution = 1.5                  (given)

volume of sol = 100/1.5  = 66.66ml  = 0.066L                          (dencity = mass/vol)

in 40gm NaOH no of moles = 40/M.wt = 40/40 = 1

molarity of sol = 1/0.066 = 15M/L

now we have M2V1 = M2V2

  15V = 0.45*500

   V =  15ml     ans

approve if u like my ans